# Efficient Ensemble Conditional Independence Test Framework for Causal Discovery

### v1

@@ Appendix F Limitations and Future Directions
The primary goal of E-CIT is a general-purpose framework for reducing the computational cost of CIT. Many in-depth discussions that require access to the alternative hypothesis distribution of the base CIT methods fall outside the main scope of this work, but they represent important directions for future research. We discuss several key points below:
Type I Error Control. Controlling Type I error is a central problem in hypothesis testing and remains one of the major challenges for CIT. In Theorem 1, we provide theoretical guarantees for Type I error control of E-CIT, in the ideal case where the subtest p-values are perfect. Although the asymptotic validity of specific CIT methods is not the focus of this work, in practice, deviations of subtest p-values from the $\text{Uniform}(0,1)$ may induce shifts in Type I error, which can be amplified by E-CIT. As illustrated in Figure 8, the p-value distributions of the original and ensemble versions of KCIT and LPCIT under the null and alternative hypotheses show that under the null, KCIT p-values are slightly deflated at both ends, while LPCIT p-values are slightly inflated. These deviations lead to a right-skew and left-skew in their ensemble versions, which may result in slightly higher or lower Type I errors.
@@
Correlated p-values. One may consider using resampling methods such as bootstrap to generate additional p-values and improve small-sample test performance. In fact, a similar idea has already been validated in small-sample causal discovery scenarios (Xiang et al., 2024). However, this approach inevitably introduces correlations among p-values. In CIT, the theoretical form of the p-value distribution under the alternative is inherently challenging, making it difficult to model correlations between tests, unlike in the parametric setting studied by (Liu & Xie, 2020; Ling & Rho, 2022). While E-CIT focuses on reducing computational cost in large-sample settings, exploring strategies to exploit overlapping data splits or correlated subtests is an interesting direction for improving CIT performance.

### v2

@@ Appendix F Limitations and Future Directions
The primary goal of E-CIT is to provide a general-purpose framework for reducing the computational cost of CIT. Many in-depth discussions that require access to the alternative hypothesis distribution of the base CIT methods fall outside the main scope of this work, but they represent important directions for future research. We discuss several key points below:
Type I Error Control. Controlling Type I error is a central problem in hypothesis testing and remains one of the major challenges for CIT. In Theorem 1, we provide theoretical guarantees for Type I error control of E-CIT, in the ideal case where the subtest p-values are perfect. Although the asymptotic validity of specific CIT methods is not the focus of this work, in practice, deviations of subtest p-values from the $\text{Uniform}(0,1)$ may induce shifts in Type I error, which can be amplified by E-CIT. As illustrated in Figure 8, the p-value distributions of the original and ensemble versions of KCIT and LPCIT under the null and alternative hypotheses show that under the null, KCIT p-values are slightly deflated at both ends, while LPCIT p-values are slightly inflated. These deviations lead to right-skew and left-skew in their ensemble versions, which may result in slightly higher or lower Type I errors.
@@
The Curse of Dimensionality. The curse of dimensionality is a fundamental challenge in CIT. While our experiments on real-world datasets encompass varying dimensions for the conditioning set $Z$, with a dedicated empirical analysis in Appendix E.3, addressing this issue directly is not the primary objective of the E-CIT framework. Instead, E-CIT is designed to be orthogonal to the internal mechanisms of base tests. Indeed, numerous specialized CIT methods (Sen et al., 2017; Bellot and van der Schaar, 2019; Shi et al., 2021; Li et al., 2023a; b) have already been developed to handle the high-dimensional conditioning sets $Z$, and E-CIT can serve as a scalable wrapper for these methods.
Correlated p-values. One may consider using resampling methods such as bootstrap to generate additional p-values and improve small-sample test performance. In fact, a similar idea has already been validated in small-sample causal discovery scenarios (Xiang et al., 2024). However, this approach inevitably introduces correlations among p-values. In CIT, the theoretical form of the p-value distribution under the alternative is inherently challenging, making it difficult to model correlations between tests, unlike in the parametric setting studied by (Liu and Xie, 2020; Ling and Rho, 2022). While E-CIT focuses on reducing computational cost in large-sample settings, exploring strategies to exploit overlapping data splits or correlated subtests is an interesting direction for improving CIT performance.
Distribution Drifts. The current theoretical guarantees of E-CIT rely on the assumption that the $K$ subtests are independent and identical, resulting in i.i.d. p-values. However, in many real-world scenarios, such as data collected across heterogeneous environments or time-series data, the underlying data-generating mechanisms may experience distribution shifts. In such instances, the theoretical guarantees of our current framework may not strictly hold. Understanding how these drifts impact the aggregated p-value, and developing adaptive subset partitioning or robust aggregation strategies to enhance resilience, represent a highly practical direction for future research.